# GeoZero: Incentivizing Reasoning from Scratch on Geospatial Scenes

### v2

@@ Appendix I Further Investigation of Reasoning Behavior
Nevertheless, it is worth emphasizing that even under the most challenging conditions, the average normalized length score $l_{s}(T)$ remains above 0.97, the average non-redundancy ratio 1d(T)1-d(T) exceeds 0.86, the average structural thinking score $q_{t}$ is greater than 0.82, and the average semantic diversity $b_{d}$ stays within a narrow range of 0.33–0.35 with only minor variation, while the average thinking-quality score $s_{t}$ also remains above 0.66. These results demonstrate that the reasoning quality across the vast majority of samples in all accuracy intervals still meets the expected standards, consistent with the findings in Figure 9. Overall, these results suggest that even as task difficulty increases, GeoZero maintains stable and meaningful reasoning performance.
@@ 5. Over what timeframe was the data collected?
K.4 Preprocessing/Cleaning/Labeling

### v3

@@ Appendix I Further Investigation of Reasoning Behavior
Nevertheless, it is worth emphasizing that even under the most challenging conditions, the average normalized length score $l_{s}(T)$ remains above 0.97, the average non-redundancy ratio 1d(T)1-d(T) exceeds 0.86, the average structural thinking score $q_{t}$ is greater than 0.82, and the average semantic diversity $b_{d}$ stays within a narrow range of 0.33–0.35 with only minor variation, while the average thinking-quality score $s_{t}$ also remains above 0.66. These results demonstrate that the reasoning quality across the vast majority of samples in all accuracy intervals still meets the expected standards, consistent with the findings in Figure 20(a). Overall, these results suggest that even as task difficulty increases, GeoZero maintains stable and meaningful reasoning performance.
@@ 5. Over what timeframe was the data collected?
K-D Preprocessing/Cleaning/Labeling